ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

13. Combinatorics

1991 MS Classification: 05A10, 05A19 Wang, Zeke, Lingnan College, Zhongshan University, P. R. China, Wang, Ergang A New Identity On Binomial Coefficients A new formulation of binomial coefficients is presented in this note to explore their alternatively-increasing-multiplier property.